OpenStudy (anonymous):

A rectangule has an area of 15 square meters. It is divided into parts each with an area of 3/8 squater meter. Into how many parts has the rectangle been divided? Help

zepdrix (zepdrix):

The number 15 is being split up into pieces of size 3/8. So you need to figure out how many 3/8's are contained in 15 wholes. Or if you prefer, you can do the `Keep Change Flip Method`\[\large\rm 15\div\frac{3}{8}\quad=\quad \frac{15}{1}\div\frac{3}{8}\quad=\quad \frac{15}{1}\times\frac{8}{3}\quad=\quad \frac{15\times8}{1\times3}\]

OpenStudy (anonymous):

Omg this is a clear answer, thank u

zepdrix (zepdrix):

cool! ୧ʕ•̀ᴥ•́ʔ୨